OpenStudy (anonymous):

Given the following matrices A and B, find an invertible matrix U such that UA = B: * matricies A and B are in comment box* I need an explanation since I am not getting this right

OpenStudy (anonymous):

\[A = \left[\begin{matrix}-4&4&-10 \\ -3 & -6 & -5\\ -1 & 2 & 5 \end{matrix}\right]\] \[B = \left[\begin{matrix}-21 & 22 & -45 \\ -3 & -6 & -5 \\ -4 & 4 & -10\end{matrix}\right]\]

OpenStudy (primeralph):



OpenStudy (anonymous):

Dont understand it, can you please go into detail?

OpenStudy (anonymous):

@primeralph please and thanks

OpenStudy (primeralph):

Can you invert a matrix?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

i'll need some guidance in going step by step. @primeralph